OpenStudy (anonymous):

f 2x + 3y = 17 and 3x + 4y = 24, then 5x +7y = 7 41 3 -7 -41

OpenStudy (anonymous):

If*

OpenStudy (shinalcantara):

for the system of equations, first thing to do is to find the values for x and y 2x + 3y = 17 ---equation 1 3x + 4y = 24 ----equation 2 can you do elimination or substitution method?

OpenStudy (anonymous):

i figured it out can you help me with another one? @shinalcantara

OpenStudy (shinalcantara):

yeah sure

OpenStudy (anonymous):

What does 1362-2(136)(36)+362=? 9448 10000 10136 76552 29584